# TabSketchFM: Sketch-based Tabular Representation Learning for Data Discovery over Data Lakes

### v1

@@ Abstract.
Enterprises have a growing need to identify relevant tables in data lakes; e.g. tables that are unionable, joinable, or subsets of each other. Tabular neural models can be helpful for such data discovery tasks. In this paper, we present TabSketchFM, a neural tabular model for data discovery over data lakes. First, we propose a novel pre-training sketch-based approach to enhance the effectiveness of data discovery techniques in neural tabular models. Second, to further finetune the pretrained model for several downstream tasks, we develop LakeBench, a collection of 8 benchmarks to help with different data discovery tasks such as finding tasks that are unionable, joinable, or subsets of each other. We then show on these finetuning tasks that TabSketchFM achieves state-of-the art performance compared to existing neural models. Third, we use these finetuned models to search for tables that are unionable, joinable, or can be subsets of each other. Our results demonstrate improvements in F1 scores for search compared to state-of-the-art techniques (even up to 70% improvement in a joinable search benchmark). Finally, we show significant transfer across datasets and tasks establishing that our model can generalize across different tasks over different data lakes.
* Joint first-authors; # Work done while at IBM Research.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://anonymous.4open.science/r/tabsketchfm-6CDC/README.md.
1. Introduction
@@
Enterprises store critical data in data lakes, large repositories of tabular data, for both governance and analytics (Nargesian et al., 2019). It is essential to effectively find relevant tables (e.g., joinable, unionable, subsets) within lakes for reporting, decision-making, statistical analysis, and more (Khatiwada et al., 2022; Khatiwada et al., 2023b). For instance, Unionable table search helps to augment an existing table with new rows, and to enrich analytics over it (Nargesian et al., 2018). Joinable table search is useful to run analytics that needs data in columns from multiple tables, e.g., to identify how many customers in some region purchased a product due to an email campaign. Also, identifying subsets or supersets of a table is essential, for example, when searching for potential copies of personal data to comply with the General Data Privacy Regulation of the European Union [^1] https://gdpr-info.eu. Motivated by such use cases, we focus on the problem of identifying tables from the data lakes that can be unionable, joinable, and/or subsets of each other  (Nargesian et al., 2018; Zhu et al., 2019; Khatiwada et al., 2023a). Specifically, in this work, we propose a pre-trained tabular model that could help in enhancing the effectiveness of data discovery techniques.
@@
Central to the problem of identifying relevant tables in the data lakes is identifying similar columns in tables. However, the columns in a table could be ambiguous. For instance, a column called Age is different when it is in a table about buildings from when it is in a table about people. Furthermore, values within a column Age play a key role in determining semantics; Age of buildings in New York are unlikely to match Age of historical buildings in Egypt. Contextualizing columns could be useful in resolving such ambiguities. Language Models (LMs) are shown to be successful in contextualizing words in a sentence  (Devlin et al., 2019). Consequently, we use them to contextualize columns and use such contextualization to determine similar columns or tables.
@@
Existing pre-trained tabular models (Deng et al., 2020; Iida et al., 2021; Wang et al., 2021a; Yin et al., 2020) based on LMs (Devlin et al., 2019) do not focus on identifying relevant tables. Instead, they focus on language-style tasks such as fact checking (Wang et al., 2021a), question answering (Herzig et al., 2020), converting natural language to SQL (Yin et al., 2020), table cell content population (Deng et al., 2020), and table metadata prediction (Iida et al., 2021; Deng et al., 2020). In addition, input to these models is the actual cell values or a row because of the focus on language style tasks. However, this has two limitations for understanding relevant tables: (a) only a small number of values can be passed in as input because of severe limits in the length of input token (e.g. 512 tokens in BERT (Devlin et al., 2019)), and this potentially limits the information available to the model to understand the input table; (b) treating numerical values in the table as text tends to lose their semantics; it maybe better to pass numerical information directly as vectors directly into the model.
@@
To address these limitations on input length and numerical representation, we introduce a novel transformer-based tabular model—TabSketchFM—exploiting both table metadata and cell values that creates different sketches over them instead of linearizing cell values. Specifically, we abstract column cell values using data sketches, as has been used in the data discovery literature (e.g., MinHash sketches  (Zhu et al., 2016), and numerical sketches  (Santos et al., 2021)). Such sketches capture tabular features, and we input them into the neural models. Inputting numerical sketches bypasses feeding table cell values to the model, addressing the token limit issue, as well as the numerical representation of cell values.
@@
Following a phase of pretraining with sketches, we fine tuned the models for the data discovery tasks of union, join, and finding subsets. While some in the literature (Fan et al., 2023) have used self supervision for fine tuning, self supervision applies better for some tasks such as unions compared to say joinability. To finetune models for the data discovery tasks, we developed a suite of 8 benchmarks across the 3 tasks LakeBench that we open source for use. On the fine tuning tasks, TabSketchFM outperformed other neural methods, by as much as 55% on F1. We then used the fine tuned models for search tasks, and showed that TabSketchFM outperforms the state-of-the-art neural and traditional methods by as much as 70% on F1.
@@
  • A novel tabular pretrained model. We are the first to present a sketch-based tabular pretrained model, TabSketchFM, which contextualizes columns based on their data sketches and tabular metadata. The sketches help to better handle input token limits and numerical cell values. To achieve this, we modified the text-based transformer architecture to take in numerical and MinHash sketches as input. A purely sketch based approach though misses the opportunity to include table semantics from column values. Given that the inputs to the transformer architecture in our system are numeric vectors, it is difficult to combine these numeric inputs with tokens derived from cell values. We therefore combine the sketch based model embeddings with off the shelf embeddings of column values from existing pretrained sentence transformer models in a novel way. We show that this combination boosts performance on tasks.
  • Finetuning benchmarks for identifying relevant tables. We develop LakeBench, a collection of 8 benchmarks that can be used to build cross encoder models to identify relevant tables. Our benchmark consists of enterprise and open data tables, and tables generated using knowledge graphs. Across the suite, we have a mix of real benchmarks and synthetically derived ones, and a variety of fine tuning tasks such as binary classification, multi-class classification, and regression.
@@
  • Demonstration of TabSketchFM on search. We compared the performance of TabSketchFM’s fine tuned models against neural and non-neural baselines. In the case of join search, TabSketchFM outperformed the nearest systems geared for join search by almost 70%. In addition, the model’s performance was robust against transfer across datasets and tasks.
2. Related Work
@@
Tabular Pretrained Models. Neural models are pretrained on tabular datasets to either recover masked tokens in the table or to detect corrupt values (Deng et al., 2020; Iida et al., 2021; Wang et al., 2021a; Yin et al., 2020). Badaro et al. (2023) surveys neural models for tabular data representation and their applications. Deng et al. (2020) combined the Masked Language Model (MLM) objective from BERT (Devlin et al., 2019) with a novel Masked Entity Recovery to train TURL. Wang et al. (2021a) use MLM with novel Cell-level Cloze and Table Context Retrival for TUTA. Yin et al. (2020) used Masked Column Prediction and Cell Value Recovery for TABERT while Iida et al. (2021) repurposed ELECTRA’s objective function (Clark et al., 2020) for TABBIE. These models are finetuned and evaluated for different downstream tasks such as table metadata prediction, table content population, fact-checking, question answering, semantic parsing, and so on. For example, TABERT is evaluated on neural semantic parsing; TABBIE for column, row, and column type population tasks; and TUTA is finetuned and evaluated for cell and table type classification tasks. These downstream tasks take either a table or a table with a question as input; they do not consider the inter-table tasks such as dataset discovery, which we consider in this paper. Furthermore, they adopt serialization techniques from LMs and represent table values as sentences, whereas we represent the tables efficiently by using sketches.
@@
Dataset Discovery Methods. Different methods such as keyword search (Cafarella et al., 2008; Brickley et al., 2019) and table search (Sarma et al., 2012; Lehmberg and Bizer, 2017; Fernandez et al., 2018) have been used for the dataset discovery over data lakes. Nargesian et al. (2018) (TUS) presented finding top-k data lake tables that are unionable with a query table. They determine the column unionability using three statistical tests based on value overlap, semantic overlap, and word embedding similarity, and aggregate them to infer table unionability. In $D^{3}L$,  Bogatu et al. (2020) adopted value overlap and word embedding measures from TUS and added three additional measures to measure column unionability: numerical column distributions, column header similarity, and regular expression matching. Khatiwada et al. (2023a) introduced SANTOS which considers the relationships between the column pairs in both query and data lake tables. Fan et al. (2023) developed a contrastive-learning-based table union search technique called Starmie that captures the context of the whole table in the form of column embeddings. Recently, Hu et al. (2023) presented AutoTUS which searches for unionable tables by contextualizing entire tables with embeddings of the relationship between the column pairs. Furthermore, there are other table search techniques that focus on finding joinable tables, i.e., given a query table marked with a query column, find the data lake tables that contain a column with overlapping values with the query column. For instance, Zhu et al. (2016) defines joinable table search as a problem of finding data lake tables that have a column having high set containment with the query column. They define a novel LSH Ensemble index that approximates set containment computation. Zhu et al. (2019) developed JOSIE that searches for top-k joinable tables based on exact set containment. Dong et al. (2021) presented PEXESO, which efficiently searches for the top-k semantically joinable tables. Along with a join on exact values, they also consider fuzzy join, i.e., join on abbreviations and synonyms. Dong et al. (2023) introduced DeepJoin, a deep-learning-based technique that also searches for semantically joinable tables. Recently, Cong et al. (2023) also proposed WarpGate, a system for searching for top-k semantically joinable tables by embedding each column to vector space using pre-trained embeddings. They index column embeddings using Locally Sensitive Hashing (Gionis et al., 1999) for efficient search. In line with these data discovery methods, we also address the task of discovering tables from the data lakes that can be unionable, joinable, and/or subsets of each other. Notably, we propose a single pretrained neural model that can be specialized for any of the above tasks. The advantage of our approach over task specific deep learning based approaches (Fan et al., 2023; Dong et al., 2023) is that, as in natural language processing, much smaller datasets are sufficient to finetune new data discovery tasks leveraging the existing tabular knowledge in the pretrained model.
@@
Benchmarks. Several tabular benchmarks exist in the literature  (Badaro et al., 2023; Hulsebos et al., 2023). Koutras et al. (2021) created benchmarks for schema matching (Rahm and Bernstein, 2001) using open data tables and evaluated the existing matching techniques over them. Mudgal et al. (2018) open-sourced 13 datasets such as the DBLP-Google Scholar author dataset and Walmart-Amazon product datasets for entity matching task (Konda et al., 2016). The WDC Web Table corpus (Lehmberg et al., 2016), containing around 233 million web tables extracted from different web pages, has been widely used for diverse tasks such as Question Answering (Herzig et al., 2020; Liu et al., 2022), Semantic Parsing (Yu et al., 2021; Liu et al., 2022; Yin et al., 2020), Table Retrieval (Wang et al., 2021b), Table Metadata Prediction (Wang et al., 2021a; Deng et al., 2020; Suhara et al., 2022) and more (Zhang and Balog, 2020). Recently, Efthymiou et al. (2022) used knowledge graphs to create tabular benchmarks for column type prediction and column-to-column binary relationship prediction tasks. The VizNet Benchmark (Hu et al., 2019) has been used to evaluate column type prediction tasks (Zhang et al., 2020; Hulsebos et al., 2019). But the most closely related benchmarks to our work are TUS (Nargesian et al., 2018), SANTOS (Khatiwada et al., 2023a), and UGEN-V1 (Pal et al., 2023) released for table union search. These search benchmarks are, however, not designed for training neural models. To create such benchmarks, we adapt the existing public datasets for pairwise classification and regression tasks. GitTables (Hulsebos et al., 2023) is a recent large table corpus in the literature. However, they contain web tables that are small entity tables with fewer rows and columns than enterprise tables and do not represent enterprise data lake scenarios (Zhu et al., 2019). We use open and enterprise datasets to create our benchmarks where the tables have 10s of columns and 1000s of rows and their cell values are often numbers or cryptic codes.
3. TabSketchFM
Figure 1. Incorporation of sketches into transformer architecture. MinHashes are computed over the whole cell value, and tokens of a cell value for a given column. MinHashes are also computed over the entire row, treated as a string. For a given column token (e.g., Price in this figure), hidden states for each encoding are summed after passing through an embedding or a linear layer before it gets sent to the BERT encoder layers. Table description tokens are summed with MinHashes over the entire row.
In this section, we detail the architecture of TabSketchFM. Like existing tabular models (Wang et al., 2021a; Herzig et al., 2020), we pretrain a transformer-based BERT model (Devlin et al., 2019), but change the architecture to accommodate numerical inputs. The goal of building the pretrained model is to create cross encoders that can be used for ranking in data discovery using significantly smaller sized labeled datasets.
The first significant departure from other tabular representation models is in how we represent the table as input. While other models linearize table cells as text, we create a sketch for each column which are numerical vectors that need to be translated into a form that can be consumed by the transformer architecture. The sketches capture different features of the columns. Figure 1 shows an overview of how the different sketches we describe below are fed to the self-attention layers. Specifically, we create the following sketches for each column.

### v3

# Harsha Kokel IBM Research

@@ Abstract
Enterprises have a growing need to identify relevant tables in data lakes; e.g. tables that are unionable, joinable, or subsets of each other. Tabular neural models can be helpful for such data discovery tasks. In this paper, we present TabSketchFM, a neural tabular model for data discovery over data lakes. First, we propose novel pre-training: a sketch-based approach to enhance the effectiveness of data discovery in neural tabular models. Second, we finetune the pretrained model for identifying unionable, joinable, and subset table pairs and show significant improvement over previous tabular neural models. Third, we present a detailed ablation study to highlight which sketches are crucial for which tasks. Fourth, we use these finetuned models to perform table search; i.e., given a query table, find other tables in a corpus that are unionable, joinable, or that are subsets of the query. Our results demonstrate significant improvements in F1 scores for search compared to state-of-the-art techniques. Finally, we show significant transfer across datasets and tasks establishing that our model can generalize across different tasks and over different data lakes.
I Introduction
@@
Enterprises store critical data in data lakes, large repositories of tabular data, for both governance and analytics [1]. It is essential to effectively find relevant tables (e.g., joinable, unionable, subsets) within lakes for reporting, decision-making, statistical analysis, and more [2, 3]. For instance, Unionable table search helps to augment an existing table with new rows, and to enrich analytics over it [4]. Joinable table search is useful to run analytics that needs data in columns from multiple tables, e.g., to identify how many customers in some region purchased a product due to an email campaign. Also, identifying subsets or supersets of a table is essential, for example, when searching for potential copies of personal data to comply with the General Data Privacy Regulation of the European Union [^1] https://gdpr-info.eu. Motivated by such use cases, we focus on the problem of identifying tables from the data lakes that can be unionable, joinable, and/or subsets of each other  [4, 5, 6]. Specifically, in this work, we propose a pre-trained tabular model that could help in enhancing the effectiveness of data discovery techniques.
@@
Central to the problem of identifying relevant tables in the data lakes is identifying similar columns in tables. However, the columns in a table could be ambiguous. For instance, a column called Age is different when it is in a table about buildings from when it is in a table about people. Furthermore, values within a column Age play a key role in determining semantics; Age of buildings in New York are unlikely to match Age of historical buildings in Egypt. Contextualizing columns could be useful in resolving such ambiguities. Language Models (LMs) are shown to be successful in contextualizing words in a sentence  [7]. Consequently, we use them to contextualize columns and use such contextualization to determine similar columns or tables.
@@
Existing pre-trained tabular models [8, 9, 10, 11] based on LMs [7] do not focus on identifying relevant tables. Instead, they focus on language-style tasks such as fact checking [10], question answering [12], converting natural language to SQL [11], table cell content population [8], and table metadata prediction [9, 8]. In addition, input to these models is the actual cell values or a row because of the focus on language-style tasks. However, this has two limitations for understanding relevant tables: (a) only a small number of values can be passed in as input because of severe limits on token length in encoder-only models that can generate embeddings for search (e.g. 512 tokens in BERT [7]), and this potentially limits the information available to the model to understand the input table 222 Decoder only models have larger context lengths. However, it has been difficult until recently to use decoder-only models for embeddings [13]. Moreover, these models are expensive to deploy and provide only a minimal improvement in accuracy over encoder-only models such as BERT. This is why embeddings for search are mostly based on BERT-style models. Further, even with large context lengths, it is not practical to serialize tables with millions of rows as sentences. ; (b) treating numerical values in the table as text tends to lose their semantics; it may be better to pass numerical information directly as vectors into the model.
@@
To address these limitations on input length and numerical representation, we introduce a novel transformer-based tabular model—TabSketchFM—exploiting both table metadata and cell values that creates different sketches over them instead of linearizing cell values. Specifically, we abstract column cell values using data sketches, as has been used in the data discovery literature (e.g., MinHash sketches  [14], and numerical sketches  [15]). Such sketches capture tabular features, and we input them into the neural models. Inputting numerical sketches bypasses feeding table cell values to the model, addressing the token limit issue, as well as the numerical representation of cell values.
@@
After pretraining using sketches, we finetune the models for individual data discovery tasks such as union, join, and finding subsets. The embeddings given by the finetuned models are then used for the respective retrieval tasks. While some in the literature [16] have used self supervision for fine tuning, self supervision applies better for some tasks such as unions compared to say joinability. To finetune models for the data discovery tasks, we used a suite of 8 benchmarks across the 3 tasks LakeBench that we have open sourced for use [17]. On the fine tuning tasks, TabSketchFM outperformed other neural methods, by as much as 55% on F1. We then used the fine tuned models for search tasks, and showed that TabSketchFM outperforms the state-of-the-art neural and traditional methods by as much as 70% on F1.
Importantly, we also compare the performance of specialized traditional and neural models for data discovery against an off-the-shelf pretrained model that encodes column values as a single sentence. To our surprise, the off-the-shelf model performs much better than existing systems - in fact, for certain tasks such as union, it is sufficient to use that model for discovering unionable tables; our tabular specific model does worse, as do all other models that target unionability. The insight from this finding is that while some data discovery tasks need values to match across columns (e.g. join or subsets), for union, it is sufficient when columns have the same semantic meaning; indeed there is no need for any value overlap. Sentence embeddings capture those semantics well. Because certain tasks seem to benefit from these sentence embeddings (e.g. union), and other do not (e.g., subset), we propose a way to concatenate the embeddings from sentence encoders with tabular model embeddings, and show that one can augment tabular embeddings with value embeddings to improve performance, at least in some cases.
@@
  • A novel tabular pretrained model. We are the first to present a sketch-based tabular pretrained model, TabSketchFM, which contextualizes columns based on their data sketches and tabular metadata. The sketches help to better handle input token limits and numerical cell values. To achieve this, we modified the text-based transformer architecture to take in numerical and MinHash sketches as input. A purely sketch based approach though misses the opportunity to include table semantics from column values. Given that the inputs to the transformer architecture in our system are numeric vectors, it is difficult to combine these numeric inputs with tokens derived from cell values. We therefore combine the sketch based model embeddings with column value embeddings from existing pretrained sentence transformer models. We show that this combination boosts performance on some data discovery tasks.
  • Finetuning models for downstream tasks. We develop cross-encoders from pretrained TabSketchFM, which we fine-tune for a range of downstream tasks from the LakeBench collection [17]. We fine tune not only TabSketchFM but other tabular pretrained models to see if they can perform well on these downstream tasks and present the comparision.
@@
  • Demonstration of TabSketchFM on search. We compared the performance of TabSketchFM’s fine tuned models against neural and non-neural baselines. In the case of join search, TabSketchFM outperformed the nearest systems geared for join search by almost 70% in F1 scores, but tellingly, the sentence encoding model outperformed existing systems as well by 63%. By combining the two embeddings (from TabSketchFM and sentence encodings), we gain almost 73% over other systems in F1.
  • Illustrate generalization. Finally, we illustrate the main advantage of learning a pretrained foundational model by showing the generalization of search performance across tasks. For example, we fine tune the pretrained model on join identification on one dataset and use it for union search on another dataset. Our evaluation shows that our finetuned models generalize well.
II Related Work
@@
Tabular Pretrained Models. Neural models are pretrained on tabular datasets to either recover masked tokens in the table or to detect corrupt values [8, 9, 10, 11]. [18] surveys neural models for tabular data representation and their applications. [8] combined the Masked Language Model (MLM) objective from BERT [7] with a novel Masked Entity Recovery to train TURL. [10] use MLM with novel Cell-level Cloze and Table Context Retrival for TUTA. [11] used Masked Column Prediction and Cell Value Recovery for TABERT while [9] repurposed ELECTRA’s objective function [19] for TABBIE. These models are finetuned and evaluated for different downstream tasks such as table metadata prediction, table content population, fact-checking, question answering, semantic parsing, and so on. For example, TABERT is evaluated on neural semantic parsing; TABBIE for column, row, and column type population tasks; and TUTA is finetuned and evaluated for cell and table type classification tasks. These downstream tasks take either a table or a table with a question as input; they do not consider the inter-table tasks such as dataset discovery, which we consider in this paper. Furthermore, they adopt serialization techniques from LMs and represent table values as sentences, whereas we represent using sketches.
@@
Dataset Discovery Methods. Different methods such as keyword search [20, 21] and table search [22, 23, 24] have been used for the dataset discovery over data lakes. Table Union Search (TUS) [4] presented finding top-k data lake tables that are unionable with a query table. They determine the column unionability using three statistical tests based on value overlap, semantic overlap, and word embedding similarity, and aggregate them to infer table unionability. $D^{3}L$ [25], adopted value overlap and word embedding measures from TUS and added three additional measures to measure column unionability: numerical column distributions, column header similarity, and regular expression matching. SANTOS[6] considers the relationships between the column pairs in both query and data lake tables. Fan et el. [16] developed a contrastive-learning-based table union search technique called Starmie that captures the context of the whole table in the form of column embeddings. AutoTUS [26] searches for unionable tables by contextualizing entire tables with embeddings of the relationship between the column pairs. Furthermore, there are other table search techniques that focus on finding joinable tables, i.e., given a query table marked with a query column, find the data lake tables that contain a column with overlapping values with the query column. For instance, LSH Ensemble [14] defines joinable table search as a problem of finding data lake tables that have a column having high set containment with the query column. They propose a novel index that approximates set containment computation. JOSIE [5] searches for top-k joinable tables based on exact set containment. AutoJoin [27], on the other hand, determines the fuzzy join between rows of a given pair of joinable tables by employing string transformations such as matching n-grams of cell values and concatenating cell values. PEXESO [28] efficiently searches for the top-k semantically joinable tables. Along with a join on exact values, they also consider fuzzy join, i.e., join on abbreviations and synonyms. DeepJoin [29] is a deep-learning-based technique that also searches for semantically joinable tables. WarpGate [30] is also a system for searching for top-k semantically joinable tables by embedding each column to vector space using pre-trained embeddings. They index column embeddings using Locally Sensitive Hashing [31] for efficient search. Recently, Chorus [32] even demonstrated the ability of Large Language Models to identify joinable columns as a task of completing ‘pd.merge‘ command when two tables are presented as pandas dataframe. Our work addresses the task of discovering tables from the data lakes that can be unionable, joinable, and/or subsets of each other. Notably, we specialize data discovery neural models based on a single pretrained model. The advantage of our approach over above mentioned task specific approaches is that, as in natural language processing, much smaller datasets are sufficient to finetune the model for new data discovery tasks; by leveraging the existing tabular knowledge in the pretrained model.
@@
Benchmarks. Several tabular benchmarks exist in the literature  [18, 33, 34]. Koutras et al. [35] created benchmarks for schema matching [36] using open data tables and evaluated the existing matching techniques over them. Mudgal et al [37] open-sourced 13 datasets such as the DBLP-Google Scholar author dataset and Walmart-Amazon product datasets for entity matching task [38]. The WDC Web Table corpus [39], containing around 233 million web tables extracted from different web pages, has been widely used for diverse tasks such as Question Answering [12, 40], Semantic Parsing [41, 40, 11], Table Retrieval [42], Table Metadata Prediction [10, 8, 43] and more [44]. Efthymiou et al. [45] used knowledge graphs to create tabular benchmarks for column type prediction and column-to-column binary relationship prediction tasks. The VizNet Benchmark [46] has been used to evaluate column type prediction tasks [47, 48]. GitTables [33] is a large table corpus in the literature. However, they contain web tables that are small entity tables with fewer rows and columns than enterprise tables and do not represent enterprise data lake scenarios [5]. But the most closely related benchmark to our work is the LakeBench collection of datasets [17] for identifying relation between pair of tables (Sec. III-D) Hence, we use that benchmark for finetuning our model. Further, there are publicly available table search benchmarks that are also relevant to our work which we use for illustrating the use of TabSketchFM for data discovery. TUS [4] and SANTOS [6] are table union search benchmarks which we include in our experiments. At the time of writing this paper, there were no search benchmarks for table join search and table subset search, so we generated Wiki join search and Eurostat subset search benchmarks (Sec. IV-C). However, recently two new benchmarks have appeared in this space. One is manually annotated table join and union search benchmark [49] (also called LakeBench [^3] Not to confuse with Srinivas et al. [17] LakeBench used for finetuning.) and another is a semantic table search [50] benchmark that used wikipedia categories and navigational links to establish semantic join and union relation.
III TabSketchFM
Figure 1: TabSketchFM Table Representation. The left panel depicts the preparation of various sketches for each column of the input table, "Residential Properties". The right panel illustrates the constituent components of the input embedding, which are subsequently summed to form the input embedding for the 12-layered BERT encoder with self-attention mechanism.
In this section, we detail the architecture of TabSketchFM. Like existing tabular models [10, 12], we pretrain the BERT model [7]. But, unlike them, we adapt the BERT input embeddings to accommodate the numerical inputs.
III-A Sketches Preparation
While other models linearize table cells as text, we prepare sketches: numerical vectors that capture different features of the table. We use MinHash signature to convert sets of strings to numerical vectors. MinHash is a technique that uses hash functions to map similar sets to similar hash values, allowing for fast estimation of Jaccard similarity between two sets [51]. [^4] We use datasketch library [52] to compute MinHash. Specifically, we create one table-level sketch called a content snapshot and two types of column-level sketches: numerical sketch and minhash sketches.
Content Snapshot. Recognizing that row information could be crucial in detecting similarity of tables, we create a sketch from the first 10000 rows. We convert each row into a string and generate a MinHash signature from the set of rows. This is the only table-level sketch that we create.